OpenStudy (anonymous):

Find the volume of the cone with the same base and height of the given cylinder.

OpenStudy (anonymous):

A. 3231.06 cm^3 B. 2236.05 cm^3 C. 1078 cm^3 D.359 cm^3

OpenStudy (accessdenied):

The volume of a cone is 1/3 the volume of a cylinder with the same dimensions. \(V = \dfrac{1}{3} \pi r^2 h \) This is always true of those shapes that come to a point in comparison to the shapes that span straight out.

OpenStudy (anonymous):

V= 1/3*1.14*r^2*7cm ?

OpenStudy (accessdenied):

\( 1/3 * 3.1415 * \underbrace{(7)}_{radius} \; ^2 * \underbrace{21}_{height} \) I use a better approximation of pi because the answer doesn't quite round with 3.14.

OpenStudy (anonymous):

1077.5345 = 1078 .. C?

OpenStudy (accessdenied):

Yup, that's what I got as well.

OpenStudy (anonymous):

Thanks, Im going to start a new question for one more.